Stock A has a beta of 0.5, and investors expect it to return 5%. Stock B has a beta of 1.5, and investors expect it to return 9%. Use the CAPM to calculate the market risk premium and the expected rate of return on the market. (Enter your answers as a whole percent.)

a.Market Risk Premium =

b.Expected Rate of Return =
Answer #1

a)

As per capital asset pricing model, required rate of return is given by

Re = Rf + ( Rm – Rf ) x Beta

Where,

Re = Required rate of return

Rf = Risk free rate of return

Rm – Rf = Market risk premium

Beta = Beta of the stock

For Stock A,

5 = Rf + (Rm – Rf) x 0.50 --------------------- ( 1 )

For Stock B,

9 = Rf + ( Rm – Rf ) x 1.5 --------------------- ( 2 )

Deducting 1 from 2 we get

9 – 5 = Rf – Rf + Market Risk Premium x ( 1.5 – 0.50 )

So, Market Risk Premium x 1 or Market Risk Premium = 9 – 5 = 4%

b)

Putting the value of market risk premium in equation 1 we get

5 = Rf + 4 x 0.50

So, Rf = 5 – 2 = 3%

Putting the value of Rf in Rm – Rf

So, expected rate of return on market ( Rm )

Rm – 3 = 4

So, Rm = 4 + 3 = 7%